OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x^2-36/42x-7x^2=(x-6)(x+6)/7x(6-x)

OpenStudy (anonymous):

\[\frac{ x^2-36}{ 42x-7x^2 }=\frac{ (x-6)(x+6) }{ 7x(6-x)}\]

OpenStudy (anonymous):

\[\frac{(x-6)(x+6)}{7x(6-x)}=\frac{(x-6)(x+6)}{-7x(x-6)}\] now, can you see it?